Write an equation in standard form of the line that passes through the given point and has the given slope m. I would like some help on solving these equations and an explanation on how to solve them. (5,-1) m=2, (-2,6) m=1, (-7,8) m=-3, (9,-10) m=-4, (-15,-4) m=1/2
Answer by stanbon(75887)

They can all be worked the same way.
Use the form y = mx + b
You are told what x,y, and m are
solve for "b"
write the equation:
------------------------------
Example:
(5,-1) m=2,
Form: y = mx + b
-1 = 2*5 + b
b = -11
Equation is y = 2x-11
---------------------------
(-2,6) m=1
y = mx + b
6 = 1*-2 + b
b = 8
y = x + 8
===========================
(-7,8) m=-3
8 = -3*7+b
b = 29
y = 3x+29
----------------
Got it ??
Cheers,
Stan H.
================
, (9,-10) m=-4, (-15,-4) m=1/2